OpenStudy (anonymous):

help show that (1-cosX)(1+1/cosX) = tan X sin X

OpenStudy (ribhu):

after taking LCM (1-cosx) (1+cosx)/(cosx) =(1-cos^2 x)/(cosx) =sin^2 x / (cosx) = sinx. sinx/cosx = sinx. tanx = RHS Hence Proved

OpenStudy (ribhu):

@clop3195 i hope i have solved ur problem